# GenSwarm: Scalable Multi-Robot Code-Policy Generation and Deployment via Language Models

### v1

@@ Abstract
The development of control policies for multi-robot systems traditionally follows a complex and labor-intensive process, often lacking the flexibility to adapt to dynamic tasks. This has motivated research on methods to automatically create control policies. However, these methods require iterative processes of manually crafting and refining objective functions, thereby prolonging the development cycle. This work introduces GenSwarm, an end-to-end system that leverages large language models to automatically generate and deploy control policies for multi-robot tasks based on simple user instructions in natural language. As a multi-language-agent system, GenSwarm achieves zero-shot learning, enabling rapid adaptation to altered or unseen tasks. The white-box nature of the code policies ensures strong reproducibility and interpretability. With its scalable software and hardware architectures, GenSwarm supports efficient policy deployment on both simulated and real-world multi-robot systems, realizing an instruction-to-execution end-to-end functionality that could prove valuable for robotics specialists and non-specialists alike.The code of the proposed GenSwarm system is available online: https://github.com/WindyLab/GenSwarm.
@@ Introduction
Automatic generation and deployment of control policies for multi-robot systems is an appealing paradigm, as it promises substantial savings in terms of human effort and other resources [3, 4, 5]. However, this paradigm is nontrivial to realize as a multi-robot system as a whole cannot be programmed directly; rather, a desired collective behavior can be achieved only by programming each individual robot, which relies on its locally available information. Previous methods for automatic development of multi-robot swarming are primarily based on optimization techniques [3, 5]. For instance, an objective function is first crafted to mathematically describe a desired task and then optimized to generate policies through methods such as evolutionary computation [6, 7, 5] or systematic search [8]. Despite their promise, these optimization methods face the common limitation of requiring manual crafting of objective functions.
@@
Recent advances in large language models (LLMs) [9, 10] and vision language models (VLMs) [11, 12] offer new paradigms for developing robotic systems. In one paradigm, a language model can be deployed onboard a robot to directly make decisions online [13, 14]. Due to the generality of language models, this paradigm could be used to address open-ended tasks [15, 16, 17]. However, it faces challenges in terms of reproducibility, interpretability, and hallucination. In another paradigm, a language model is used to generate executable code policies that are subsequently uploaded for execution on-board robots. A representative method that falls into this paradigm is Code-as-Policy [18, 19, 20]. Due to the white-box nature of executable code, this paradigm offers high reproducibility and interpretability. Moreover, since executable code usually requires fewer resources than LLMs, this paradigm also enables real-time control on low-cost robot platforms. This is especially relevant for large-scale swarming systems, where collective behaviors emerge from the interactions among robots with exceedingly limited onboard resources [21, 22, 23, 24, 25, 26]. Therefore, this code-policy paradigm is adopted in the present work.
@@
Despite the promise of the code-policy paradigm, current studies mainly focus on single-robot systems [18, 27, 28]. For multi-robot systems, the development of control policies faces additional challenges. First, the design of policies must consider a robot’s interactions with its peers. In some situations the robot may compete with its peers, for example, for limited resources, whereas in others it may cooperate with its peers to achieve a common goal [26, 29, 5]. Second, the deployment and maintenance of policies require scalable software and hardware systems, which is particularly relevant for robotic swarms that may have a large number of robots. Third, to maximize the utility of a multi-robot system, it needs to support a wide range of tasks. Although some studies used LLMs to realize specific tasks, such as cooperative navigation [30], flocking [31], dancing [32, 33], and manipulation [34], the generality of the proposed methods, beyond the specific use case, is yet to be demonstrated. A few further studies have explored LLMs with groups of agents [35, 36, 37], but the agents therein are simulated non-embodied agents; hence the findings may be difficult to translate to real multi-robot systems.
Here, we propose GenSwarm, an end-to-end system that can automatically generate and deploy code policies for versatile multi-robot tasks. GenSwarm enables users to program a group of robots using simple natural language instructions. The user instructions are automatically processed via a pipeline of components, including constraint analysis, policy design, policy generation, policy deployment in simulation environments, policy deployment on real-world robots, and policy improvement based on feedback. These components are respectively empowered by LLM agents. Therefore, GenSwarm realizes generative multi-robot swarming via the cooperation among multiple language agents.
@@
GenSwarm can deploy the generated code policies as well as the required runtime environments fully automatically on simulated and real-world robots, thus achieving true end-to-end functionality. The deployment is facilitated by a scalable multi-robot platform that features novel software and hardware architectures. GenSwarm enables zero-shot policy generation without the need for context learning based on demonstrative examples. When altered or unseen tasks arise, the system can re-generate and re-deploy policies in response to user requests, thereby offering high adaptability for dynamic tasks. Furthermore, due to the use of code policies, the approach is not only reproducible and interpretable, but also suitable for real-time execution on robots with limited onboard resources.
@@
The code deployment and improvement module realizes automatic code deployment in simulated and real-world robotic platforms (Fig. 1c). It relies on novel hardware and software systems, which will be detailed in the following section. It introduces multi-modal feedback mechanisms that can automatically identify issues during execution and effectively adjust policies based on feedback. Specifically, execution results in the simulation can be automatically collected in the format of video clips. A VLM agent assesses the video clips to generate feedback on whether the desired task is successfully completed. In addition, an interface for human feedback is incorporated to enable users to efficiently modify the policy by providing natural language feedback.
@@
Automatic deployment is nontrivial as code execution depends on complex runtime environments consisting of various software packages. Manually installing and configuring the runtime environments on each robot would be inefficient as the time required scales linearly with the number of robots; for example, deploying on 10 robots would take 10 times longer than deploying on one robot. This would make deployment on large-scale swarms impractical.
@@
GenSwarm possesses a software framework that can automatically deploy both the runtime environments and the generated code across all the robots in near constant time regardless of the number of robots. In our experiments, automatically deploying the runtime environments on all the robots takes about two minutes, whereas automatically deploying the generated code takes mere seconds. This makes the system particularly well-suited for large-scale swarming systems, where consistent and rapid deployment is essential.
@@
The software framework is illustrated in Fig. 2. A control station first generates the required code based on the pipeline described earlier and connects with each robot through Ansible via WiFi and SSH (Methods). With predefined automated scripts in the format of Playbook, each robot performs a series of tasks such as installing and configuring the Docker environment. After the Docker environment is ready, two pre-built Docker images are pulled: one containing the ROS environment used for robot operation, and the other containing the Python environment required for code execution. Once the execution environments are ready, the generated code is transmitted to all robots and then executed onboard. The proposed software framework heavily relies on two techniques, Ansible and Docker (Methods), which work together to simplify and streamline the code deployment on multiple robots. This integration ensures that the deployment process is both repeatable and efficient, drastically reducing the time required to make a group of robots operational. In addition, the framework is realized by a modular architecture, whose details are given in Methods.
The hardware framework is illustrated in Fig. 2. A new multi-robot platform, which is a major upgrade of our previous robotic platform [26], was developed to support GenSwarm. Each ground robot has onboard computational, control, and communication resources that are necessary for autonomous code deployment and execution [38]. Considering that multi-robot experiments involve a large number of operations, such as starting and shutting down robots, we developed novel features for the multi-robot platform such as one-click all start, one-click all sleep, and wireless data retrieval, significantly reducing experimental costs.
It is worth mentioning that since the robots do not possess onboard vision systems, the perception is emulated in a distributed manner. Specifically, relevant motion information of all the robots is collected by an indoor positioning system first and then distributed to all the robots through an MQTT coordination server so that each robot receives only information about its surroundings (Fig. 3). In the future, this coordination server could be replaced with on-board vision systems. Another merit of the MQTT coordination server is to enable large-scale hardware-in-the-loop swarming simulation and seamless transition from simulation to real-world implementation [38].
@@
Based on those skills and the consequently generated skill graph that describes their hierarchical dependencies, LLM agents further generate and then review the main-body code of each skill function (Fig. 4d). Logical or grammatical code errors can be identified and corrected. Once the skill functions pass the review process and static code analysis, they are automatically deployed and executed in the simulation environment. Then, a VLM agent reviews the video clip of the simulation execution and provides feedback for improving the code (Fig. 4e). After that, the generated code is automatically deployed on real-world robotic platforms. It is worth mentioning that human feedback can be incorporated to adjust the code policy (Fig. 4e). For instance, if the human feedback is “Forming a circle with a radius of 1 meter seems a bit large; perhaps 0.8 meters would be better.”, GenSwarm can adjust the corresponding parameter from 1 to 0.8, enabling efficient human-in-the-loop policy adjustment.
@@
Non-stop one-take videos are attached to show the complete workflows of GenSwarm (Movies 1 and 2). In terms of time consumption, the steps of code generation, deployment onto real-world robots, and improvement based on human feedback took approximately six, two, and two minutes, respectively. The time duration of code generation can be significantly shortened if LLMs’ efficiency can be improved in the future. The time of deployment can be shortened to a few seconds if the runtime environment has been pre-installed on the robots and merely the generated code needs to be deployed.
As elaborated above, GenSwarm consists of multiple LLM agents that play different roles. Each LLM agent is set up in advance by a prompt involving role description, environment description, robot description, and available APIs. For instance, regarding environment description, the prompt may be “The environment is composed of a 2D plane with obstacles and robots”. Regarding robot description, the prompt may be “The maximum speed of each agent is 0.5 m/s”. Regarding APIs, the prompt may be “There are two types of APIs: local and global. Local APIs can only be called by the robot itself, and global APIs can be called by a cen-
tralized controller
@@
”. Examples of local APIs are “get_self_position” and “get_surrounding_
robots_info
”. Examples of global APIs are “get_all_robots_id” and “get_all_robots_initial_
position
”. More information can be found in the code on GitHub (Code Availability).
@@
Fig. 6: Evaluation results. a. The success rates of GenSwarm across ten multi-robot tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each task. Hence, 1,000 trials in total were run and the average success rate was 81%. b. The comparison between GenSwarm, CaP, MetaGPT, and GenSwarm without VLM feedback across six representative tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. The average success rates of GenSwarm, GenSwarm without VLM, CaP, and MetaGPT were 74%, 71%, 40%, and 31%, respectively.
Performance Evaluation
The performance of GenSwarm was evaluated on ten multi-robot tasks, including aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering (Fig. 5). These tasks cover a wide range of scenarios, from cooperative to competitive, aiming to comprehensively evaluate the effectiveness of GenSwarm. Details of the tasks and the evaluation metrics are given in Methods. The LLM used here was o1-mini, one of the state-of-the-art LLMs. One hundred independent trials, starting from user instruction to code execution in simulation, were run for each of the ten tasks. The average success rate over the 1,000 trials for 10 tasks was 81%. The respective success rate for each task is presented in Fig. 6a.
The performance of GenSwarm is also compared to two state-of-the-art methods, MetaGPT [35] and Code-as-Policy (CaP) [18]. MetaGPT and CaP are two seminal methods that represent the mainstream paradigms in the fields of complex software generation and robot policy generation, respectively. GenSwarm without VLM feedback was also evaluated. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. Six representative tasks were selected, and hence 1,800 trials in total were run. The LLM used here was GPT-4o. The comparison results are shown in Fig. 6b. As can be seen, GenSwarm achieved the highest average success rate, which was 74%, across different tasks. The average success rates of GenSwarm without VLM, CaP, and MetaGPT were 71%, 40%, and 31%, respectively. It is worth mentioning that to avoid underestimating the capabilities of CaP and MetaGPT, we discarded trials that were not executable due to code bugs and only counted the executable trials for CaP and MetaGPT. Nevertheless, the two methods still show lower performance than GenSwarm. By comparing Figs. 6a and 6b, it can be seen that different LLMs (o1-mini and GPT-4o) lead to similar success rates though there are subtle variations.
The user instructions have a significant impact on the success rate of GenSwarm (Extended Data Fig. 1). The success rate tends to be higher when comprehensive instructions are provided. When the instructions are ambiguous, the generated policies are more likely to encounter issues. Note that due to the limited reasoning ability and hallucinations of LLMs, even detailed instructions cannot guarantee a 100% success rate (Extended Data Fig. 1a). Moreover, it is still possible, albeit unlikely, to successfully accomplish the task when the instructions are ambiguous (Extended Data Fig. 1b-c).

### v2

@@ Abstract
The development of control policies for multi-robot systems traditionally follows a complex and labor-intensive process, often lacking the flexibility to adapt to dynamic tasks. This has motivated research on methods to automatically create control policies. However, these methods require iterative processes of manually crafting and refining objective functions, thereby prolonging the development cycle. This work introduces GenSwarm, an end-to-end system that leverages large language models to automatically generate and deploy control policies for real-world multi-robot systems based on user instructions in natural language. As a multi-language-agent system, GenSwarm achieves zero-shot learning, enabling rapid adaptation to altered or unseen tasks. The white-box nature of the code policies ensures strong reproducibility and interpretability. With its scalable software and hardware architectures, GenSwarm supports efficient and automated policy deployment on both simulated and real-world multi-robot systems, realizing an instruction-to-execution end-to-end functionality that may transform the development paradigm of multi-robot systems in the future.
@@ Introduction
Automatic generation and deployment of control policies for multi-robot systems is an appealing paradigm, as it promises substantial savings in terms of human effort and other resources [3, 4, 5]. However, this paradigm is nontrivial to realize as a multi-robot group as a whole cannot be programmed directly; rather, a desired collective behavior can be achieved only by programming each individual robot, which relies on its locally available information. Previous methods for automatic development of multi-robot systems are primarily based on optimization techniques [3, 5]. For instance, an objective function is first crafted to mathematically describe a desired task and then optimized to generate policies through methods such as evolutionary computation [6, 7, 5] or systematic search [8]. Despite their promise, these optimization methods face the common limitation of requiring manual crafting of objective functions.
@@
Recent advances in large language models (LLMs) [9, 10] and vision language models (VLMs) [11, 12] offer new paradigms for developing robotic systems. In one paradigm, a language model can be deployed onboard a robot to directly make decisions online [13, 14, 15, 16]. Due to the generality of language models, this paradigm could be used to address open-ended tasks [17, 18, 19]. However, it faces challenges in terms of reproducibility, interpretability, and hallucination. In another paradigm, a language model is used to generate executable code policies that are subsequently uploaded for execution on-board robots. A representative method that falls into this paradigm is Code-as-Policy (CaP) [20, 21, 22]. Due to the white-box nature of executable code, this paradigm offers high reproducibility and interpretability. Moreover, since executable code usually requires fewer resources than LLMs, this paradigm also enables real-time control on low-cost robot platforms. This is especially relevant for large-scale multi-robot systems, where collective behaviors emerge from robots with exceedingly limited onboard resources [23, 24, 25, 26, 27, 28]. Therefore, this code-policy paradigm is adopted in our work.
@@
Despite the promise of the code-policy paradigm, the development of control policies for multi-robot systems faces additional challenges compared to single-robot systems [20, 29, 30]. First, the design of policies must consider a robot’s interactions with its peers. In some situations, the robot may compete with its peers, for example, for limited resources, whereas in others it may cooperate with its peers to achieve a common goal [28, 31, 5]. Second, the deployment and maintenance of policies require scalable software and hardware systems, which is particularly relevant for multi-robot systems that may have a large number of robots. Third, to maximize the utility of a multi-robot system, it needs to support a wide range of tasks. In addition, some studies proposed frameworks for automated software development such as MetaGPT [32], ChatDev [33], and [34]. Although broadly relevant, these frameworks are not specifically designed for multi-robot systems.
Recently, a number of studies explored the use of LLMs for multi-robot systems, but their applicability to general-purpose and real-world multi-robot systems still faces significant hurdles. Of particular relevance is LLM2Swarm [35], which takes user instructions as input and outputs control policies for individual robots. Although LLM2Swarm is intended to be task-agnostic, its generality is yet to be experimentally verified. Moreover, LLM2Swarm depends on manually-written demonstration examples, restricting its zero-shot capabilities. Other methods such as SmartLLM [36] focus on high-level symbolic planning and do not generate executable low-level control policies. Furthermore, many methods are tailored for specific tasks—such as formation control [37, 38], cooperative navigation [39], dancing [40, 41], or manipulation [42]—and thus lack the generality to address multiple multi-robot tasks. Moreover, the validation in most of the aforementioned methods is performed in simulation, leaving the significant challenge of automated policy deployment on physical multi-robot systems largely unexplored.
@@
Here, we propose GenSwarm, an end-to-end system that can automatically generate and deploy multi-robot policies on real-world platforms from natural language instructions for versatile multi-robot tasks. GenSwarm enables users to program a group of robots using simple natural language instructions. The user instructions are automatically processed via a pipeline of components, including constraint analysis, policy design, policy generation, policy deployment in simulation environments, policy deployment on real-world robots, and policy improvement based on feedback. These components are respectively empowered by LLM agents. GenSwarm can automatically deploy the generated code policies as well as the required runtime environments on real-world robots, thus achieving true end-to-end functionality. The automatic deployment is realized by a scalable multi-robot platform that features novel software and hardware architectures. GenSwarm enables zero-shot policy generation without the need for context learning based on demonstrative examples. When altered or unseen tasks arise, the system can re-generate and re-deploy policies in response to user requests, thereby offering high adaptability for dynamic tasks. Furthermore, due to the use of code policies, the approach is suitable for real-time execution on robots with limited onboard resources.
Extensive experiments demonstrate the high success rate of GenSwarm across various multi-robot tasks. GenSwarm consistently outperforms the state-of-the-art methods including MetaGPT [32], CaP [20], and LLM2Swarm [35], achieving significant improvements of 37%, 34%, and 34% in average success rate. GenSwarm provides a promising new paradigm for developing multi-robot systems. Its significance lies in overcoming two limitations of existing work. First, developing multi-robot systems is time-consuming and labor-intensive, and this problem worsens as the number of robots increases. Second, current multi-robot systems lack generality and flexibility. They are often limited to specific tasks or cannot adapt to changing goals and new situations in a timely manner. GenSwarm overcomes these limitations and has the potential to transform the development paradigm of multi-robot systems.
Results
@@
The code deployment and improvement module realizes automatic code deployment in simulated and real-world robotic platforms (Fig. 1c). It relies on novel hardware and software systems, which will be detailed in the following section. It introduces multi-modal feedback mechanisms that can automatically identify issues during execution and effectively adjust policies based on feedback. Specifically, execution results in the simulation can be automatically collected in the format of video clips. A VLM agent assesses the video clips to generate feedback on whether the desired task is successfully completed. In addition, an interface for human feedback is incorporated. It enables users to efficiently modify the policy by providing natural language feedback.
The global-local control structure can be automatically determined and implemented by the proposed pipeline. During the task analysis stage, the LLM agent judges whether the task requires global skills for global centralized coordination or merely local skills for local distributed execution. This structural decision is encoded in the skill graph to guide code generation, and the resulting architecture determines the deployment model: if global skills are generated, they execute once on the control station using global information, after which local skills are deployed to each robot for distributed execution based only on the local information exposed by the system APIs.
@@
Automatic deployment is nontrivial as code execution depends on complex runtime environments consisting of various software packages. Manually installing and configuring the runtime environments on each robot would be inefficient as the time required scales linearly with the number of robots. This would make deployment on large-scale multi-robot systems impractical.
@@
GenSwarm possesses a software framework that can automatically deploy both the generated code and the runtime environments across all the robots in near constant time regardless of the number of robots. In our experiments, automatically deploying the runtime environments on all the robots takes about two minutes, whereas automatically deploying the generated code takes mere seconds. This makes the system particularly well-suited for large-scale multi-robot systems, where consistent and rapid deployment is essential.
@@
The software framework is illustrated in Fig. 2. A control station first generates the required code based on the pipeline described earlier and connects with each robot through Ansible via WiFi and SSH (Methods). With predefined automated scripts in the format of Playbook, each robot performs a series of tasks such as installing and configuring the Docker environment. After the Docker environment is ready, two pre-built Docker images are pulled: one containing the ROS environment used for robot operation, and the other containing the Python environment required for code execution. Once the execution environments are ready, the generated code is transmitted to all robots and then executed onboard. The proposed software framework heavily relies on two techniques, Ansible and Docker (Methods), which work together to simplify and streamline the code deployment on multiple robots. This integration ensures that the deployment process is both repeatable and efficient, drastically reducing the time required to make a group of robots operational. Moreover, the framework is designed to be portable across different hardware platforms, a feature enabled by its modular software architecture, which is detailed in the Methods section.
The hardware framework is illustrated in Fig. 3. A new multi-robot platform, which is a major upgrade of our previous robotic platform [28], was developed to support GenSwarm. Each ground robot has onboard computational, control, and communication resources that are necessary for autonomous code deployment and execution [43]. Considering that multi-robot experiments involve a large number of operations, such as starting and shutting down robots, we developed novel features for the multi-robot platform such as one-click all start, one-click all sleep, and wireless data retrieval, significantly reducing experimental costs. It is worth mentioning that the perception of each robot is emulated. Specifically, the motion information of all the robots is collected by an indoor positioning system and then distributed to all robots through an MQTT coordination server so that each robot receives only information about its surroundings (Fig. 3). The generated code policies access the required information by calling APIs (application programming interfaces) that enforce hard-coded physical limitations. For instance, the sensing API restricts a robot’s perception to a fixed local radius (1 m in our experiments), while the motion API clamps velocity commands to a predefined maximum speed. In the future, the sensing API could be realized by onboard vision systems. As the indoor positioning system can provide high-precision measurements, we also exposed our multi-robot system to different levels of measurement noise in real-world experiments. Although performance gradually deteriorates as the noise level increases, the system is still reasonably stable for low to moderate levels of noise. The real-world noise robustness results are provided in Supplementary Fig. 1.
@@
For tasks like flocking and aggregation that do not involve global goals or constraints, GenSwarm generates distributed policies whose execution merely relies on local information. In contrast, for tasks like shaping that involve global goals or constraints, GenSwarm usually generates combinations of centralized coordination (e.g., position assignment) and distributed control. It is notable that GenSwarm automatically selects and generates control structures, ranging from pure distributed control to hybrid centralized coordination plus distributed control, which reflect the characteristics of the task. This selection process leverages the LLM’s strong prior knowledge, learned from its vast training data of robotics literature and code, to associate a high-level task description with a typical and effective control paradigm. In the resulting hybrid architecture, a global skill runs only once on the control station to perform one-time centralized coordination (e.g., initial goal assignment). After this, each robot executes local skills in a distributed manner, relying entirely on local information that is strictly enforced by the system’s sensing and motion APIs. The flexibility of augmenting pure distributed control with hybrid centralized coordination allows GenSwarm to adapt across a wide range of multi-robot tasks.
@@
Based on those skills and the consequently generated skill graph that describes their hierarchical dependencies, LLM agents further generate and then review the main-body code of each skill function (Fig. 4d). Logical or grammatical code errors can be identified and corrected. Once the skill functions pass the review process and static code analysis, they are automatically deployed and executed in the simulation environment. Then, a VLM agent reviews the video clip of the simulation execution and provides feedback for improving the code (Fig. 4e). After that, the generated code is automatically deployed on real-world robotic platforms. It is worth mentioning that human feedback can be incorporated to adjust the code policy (Fig. 4e). For instance, if the human feedback is “Forming a circle with a radius of 1 meter seems a bit large; perhaps 0.8 meters would be better.”, GenSwarm can adjust the corresponding parameter from 1 to 0.8, enabling efficient human-in-the-loop policy adjustment. The ability of human-in-the-loop adjustment provides a practical approach to adapt to newly emerged situations such as robot faults. While this adaptation may not occur in real-time (e.g., at millisecond-level latency), it still offers an effective way for rapid reprogramming and redeployment.
@@
Non-stop one-take videos are attached to show the complete workflows of GenSwarm (Movies 1 and 2). In terms of time consumption, the steps of code generation, deployment onto real-world robots, and improvement based on human feedback took approximately six, two, and two minutes, respectively. The time duration of code generation can be significantly shortened if LLMs’ efficiency can be improved in the future. The time of deployment can be shortened to a few seconds if the runtime environment has been pre-installed on the robots and merely the generated code needs to be deployed. As elaborated above, GenSwarm consists of multiple LLM agents that play different roles. All LLMs and VLMs in GenSwarm are used out-of-the-box without fine-tuning. This design was made to maximize reproducibility, enabling any user to directly deploy the system using off-the-shelf models. Each LLM agent is set up in advance by a prompt involving role description, environment description, robot description, and available APIs. For instance, regarding environment description, the prompt may be “The environment is composed of a 2D plane with obstacles and robots”. Regarding robot description, the prompt may be “The maximum speed of each agent is 0.5 m/s”. Regarding APIs, the prompt may be “There are two types of APIs: local and global. Local APIs can only be called by the robot itself, and global APIs can be
called by a centralized controller
@@
”. Examples of local APIs are “get_self_position” and “get_surrounding_robots_info”. Examples of global APIs are “get_all_robots_id” and “get_all
_robots_initial_position
”. While merely some representative examples are provided here, the complete prompts and APIs can be found in our open-source repository (see Code Availability).
Performance Evaluation
Different tasks: The performance of GenSwarm was evaluated on ten different multi-robot tasks, including aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering (Fig. 5). These tasks cover a wide range of scenarios, from cooperative to competitive, aiming to comprehensively evaluate the effectiveness of GenSwarm. Details of the tasks and the evaluation metrics are given in Methods. The LLM used here was o1-mini, one of the state-of-the-art LLMs. One hundred independent trials, starting from user instruction to code execution in simulation, were run for each of the ten tasks. The average success rate over the 1,000 trials for 10 tasks was 81%. The respective success rate for each task is presented in Fig. 6a.
@@
Fig. 6: Success rate of different LLMs on different tasks. a. The success rates of GenSwarm across ten multi-robot tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each task. Hence, 1,000 trials in total were run and the average success rate was 81%. b. The comparison between GenSwarm, CaP, MetaGPT, LLM2Swarm and GenSwarm without VLM feedback across six representative tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. The average success rates of GenSwarm, GenSwarm without VLM, CaP, MetaGPT, and LLM2Swarm were 74%, 71%, 40%, 37%, and 40%, respectively.
Different methods: GenSwarm has been compared to three state-of-the-art methods, MetaGPT [32], CaP [20], and LLM2Swarm [35]. To ensure a fair comparison, all baseline frameworks were configured according to their native design paradigms (zero-shot or few-shot). For few-shot methods like CaP and LLM2Swarm, we provided high-quality multi-robot examples—handcrafted for CaP and drawn from the official LLM2Swarm repository for the latter—with minimal adaptations for our platform. The complete prompts and code examples used for all baselines are publicly available for reproducibility (see Code Availability). Moreover, GenSwarm without VLM feedback was also compared. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. Six representative tasks were selected, and hence 2,400 trials in total were run. The LLM used here was GPT-4o. The comparison results are shown in Fig. 6b. As can be seen, GenSwarm achieved the highest average success rate, which was 74%, across different tasks. The average success rates of GenSwarm without VLM, CaP, MetaGPT, and LLM2Swarm were 71%, 40%, and 37%, and 40%, respectively. GenSwarm consistently outperforms the baselines, achieving 34%, 37%, and 34% higher success rate than LLM2Swarm, MetaGPT, and CaP. We also compared using finer-grained task-specific metrics, with detailed results presented in Supplementary Fig. 2. It reveals that GenSwarm consistently achieves superior performances across different tasks.
In addition, Supplementary Fig. 2 includes a benchmark against fine-tuned state-of-the-art (SOTA) expert controllers. For distributed tasks such as flocking and aggregation, we used the classic Boids model [45]; for hybrid tasks such as shaping, crossing, covering, and encircling, we combined an optimal assignment algorithm based on the Hungarian method with VR-ORCA [46]. While these SOTA controllers achieve higher average performance, GenSwarm’s best-performing policies reach comparable levels in some cases, demonstrating its potential to deliver high-quality solutions without extensive manual tuning.
Different LLMs: By comparing Figs. 6a and 6b, it can be seen that different LLMs (o1-mini and GPT-4o) lead to similar success rates though there are subtle variations. We further expanded the comparison to include two additional prominent LLMs, DeepSeek-V3 and Claude-3.7-Sonnet. Results consistently show high success rates across these models (Supplementary Fig. 3), suggesting general applicability of GenSwarm across different types of LLMs.
Different prompts: The user instructions have a significant impact on the performance of GenSwarm. For instance, comprehensive instructions tend to yield better results, while ambiguous ones may lead to failures (see examples in Supplementary Fig. 4). To systematically analyze this effect, we designed seven representative prompt types that range from unstructured to highly structured formats: 1) Plain-Compound (Cohesive), which integrates both the task objective and policy into a linguistically coherent paragraph; 2) Plain-Compound, which strictly concatenates the verbatim text from the objective-only and policy-only prompts; 3) Plain-Objective, which provides only the objective but no policy; 4) Plain-Policy, which provides only the policy but omits task objectives; 5) Plain-Narrative, which uses natural, human-like language to describe the task but lacks formal structure or policy details; 6) Structured-Objective, which restructures the instruction into a “description-goal-constraint” format; and 7) Structured-Policy, which adds explicit policies and constraints on top of the structured prompt. Examples of the seven prompt types for an encircling task are given in Supplementary Fig. 5. Moreover, all the previous evaluations were conducted using the Plain-Compound (Cohesive) prompt type.
As shown in Supplementary Fig. 6, the inclusion of explicit policy instructions is helpful for achieving high task success rates. Specifically, prompt types that contained policy instructions— Plain-Compound (Cohesive) (78%), Plain-Compound (74%) , Plain-Policy (74%), and Structured-Policy (74%)—yielded higher success rates. Conversely, prompts lacking this information, such as Plain-Objective (56%), Plain-Narrative (57%), and Structured-Objective (57%), resulted in significantly lower success rates. This demonstrates that the presence of policy instructions is more impactful than the prompt’s format (i.e., natural vs. structured language). Among the top performers, Plain-Policy emerges as a particularly practical choice, leading to high success rates while offering the simplicity of concise natural language inputs.